# Batch learning equals online learning in Bayesian supervised learning

### v1

@@ 5. Probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$
In this section we assume that $\mathcal{Y}$} is a Polish space unless otherwise stated. Then ${\mathcal{P}}({\mathcal{Y}})^{k}$ and ${\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{k})$ are Polish spaces for any $k\in{\mathbb{N}}^{+}$. Using Orbanz’s description of the space ${\mathcal{P}}^{2}({\mathcal{Y}})$ for a Polish space $\mathcal{Y}$} [Orbanz2011, Theorem 1.1], we shall describe the space ${\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ using a projective system (Theorem 5.5).
@@
Denote by $R^{\mathcal{X}}_{X_{m}}$ the restriction map ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{{\mathcal{X}}_{m}}$.
@@
Conversely, if $\mathcal{Y}$} is a Polish space, for any projective system of probability spaces $\{({\mathcal{P}}({\mathcal{Y}})^{X_{m}},\mu_{X_{m}}),R_{X_{m}}^{X_{n}}:X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})\}$ there exists a unique probability measure $\mu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ such that for all $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ we have
@@
Applying the Kolmogorov extension theorem, we obtain immediately Lemma 5.1. ∎
@@
Let ${\mathcal{A}}_{\mathcal{Y}}$ be the algebra consisting of open balls with rational radius centered at a countable dense set in $\mathcal{Y}$}. Then ${\mathcal{A}}_{\mathcal{Y}}$ generates the Borel $\sigma$-algebra of $\mathcal{Y}$}. Set
@@
$\widehat{e}v_{\mathcal{Y}}^{(k)}:{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{k}\big)\to{\mathcal{P}}({\mathcal{Y}})^{k}$
@@
(5.2) $\widehat{ev}_{\mathcal{Y}}^{(k)}(\nu)(B_{1}\times\ldots\times B_{k}):=\int_{{\mathcal{P}}({\mathcal{Y}})^{k}}\mu(B_{1}\times\ldots\times B_{k})\,d\nu(\mu)$
@@
$\widehat{e}v_{\mathcal{Y}}^{X_{m}}:{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{m}}\big)\to{\mathcal{P}}({\mathcal{Y}})^{X_{m}}$
@@
(5.3) $\widehat{ev}_{\mathcal{Y}}^{X_{m}}(\nu)(B_{1}\times\ldots\times B_{m}):=\int_{{\mathcal{P}}({\mathcal{Y}})^{X_{m}}}\mu(B_{1}\times\ldots\times B_{m})\,d\nu(\mu)$
@@
(5.4) $\widehat{ev}^{(1)}_{\mathcal{Y}}=P_{*}ev_{\mathcal{Y}}.$
Consequently, we have
(5.5) $\widehat{ev}_{\mathcal{Y}}^{(1)}\circ P_{*}\delta=\operatorname{Id}_{{\mathcal{P}}({\mathcal{X}})}.$
@@
(5.6) $\widehat{ev}_{\mathcal{Y}}^{\mathcal{X}}(\nu):=\lim_{\stackrel{{\scriptstyle\leftarrow}}{{X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})}}}\widehat{ev}_{\mathcal{Y}}^{X_{m}}\big((R^{\mathcal{X}}_{X_{m}})_{*}\nu\big)\in{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}},$
@@
(1) For $\nu\in{\mathcal{P}}^{2}({\mathcal{Y}})$ and $B\in\Sigma_{\mathcal{Y}}$ we compute:
$\widehat{ev}_{\mathcal{Y}}^{(1)}(\nu)(B)=\int_{{\mathcal{P}}({\mathcal{Y}})}\mu(B)d\nu(\mu)=\int_{{\mathcal{P}}({\mathcal{Y}})}\overline{ev}_{\mathcal{X}}(B|\mu)d\nu(\mu)=P_{*}ev_{\mathcal{X}}(\nu),$
which proves the equality (5.4).
To prove (5.5), by [Chentsov72, Lemma 5.10, p. 73], see also [Giry82, Theorem 1], taking into account (5.4), it suffices to show that
(5.7) $P_{*}(ev_{\mathcal{Y}}\circ\delta)=P_{*}\operatorname{Id}_{{\mathcal{X}}}.$
Recalling following formula [JLT21, (10)] for $T\in{\bf Probm}({\mathcal{X}},{\mathcal{Y}})$
$P_{*}(T)=\widehat{ev}_{\mathcal{X}}\circ P_{*}(\overline{T})$
we conclude that the RHS of (5.7) is equal $\widehat{ev}_{\mathcal{X}}\circ P_{*}(\delta)$, which is equal the LHS of (5.7) by Formulas (2.3) and (5.4).
(2) Denote by $\Pi_{i}:{\mathcal{P}}({\mathcal{Y}})^{k}\to{\mathcal{P}}({\mathcal{Y}})$ the projection onto the $i$-th component. To prove the second assertion of Lemma 5.4, it suffices to show that for any $i\in\overline{1,k}$ the composition $\Pi_{i}\circ\widehat{ev}_{\mathcal{Y}}$ is measurable. Noting that
(5.8) $\Pi_{i}\circ\widehat{ev}^{k}_{\mathcal{X}}=\widehat{ev}^{(1)}_{\mathcal{X}}$,
and using the first assertion of Lemma 5.4, we conclude that $\Pi_{i}\circ\widehat{ev}^{k}_{\mathcal{Y}}$ is measurable, what is required to prove.
Now assume that $\mathcal{Y}$} is a Polish space. The $(\tau_{w},\tau_{w})$-continuity assertion can be proved in the same way, as the measurable assertion, using [Giry82, Theorem 1]. Here we offer a slightly different argument. By [Giry82, Theorem 1] the map $\widehat{ev}_{{\mathcal{X}}^{k}}:{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}}^{k}))\to{\mathcal{P}}({\mathcal{Y}}^{k})$ is $(\tau_{w},\tau_{w})$-continuous. Now we consider the following diagram:
(5.9) ${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{k}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{ev}_{\mathcal{Y}}^{(k)}}$$\scriptstyle{P_{*}\mathfrak{m}^{k}}$${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}}^{k})\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{ev}_{{\mathcal{Y}}^{k}}}$${{\mathcal{P}}({\mathcal{Y}})^{k}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathfrak{m}^{k}}$${{\mathcal{P}}({\mathcal{Y}}^{k}).}$
One verifies immediately that the diagram is commutative. Noting that $\mathfrak{m}^{k}$, $P_{*}\mathfrak{m}^{k}$ are continuous and injective, the continuity of $\widehat{ev}_{\mathcal{Y}}^{(k)}$ follows from the continuity of the maps $P_{*}T$ and $\widehat{ev}_{{\mathcal{Y}}^{k}}$.
3) The last assertion of Lemma 5.4 follows from the second one, taking into account the Kolmogorov extension theorem and the commutativity of the following diagram for any $X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})$:
(5.10) ${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{n}}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}R^{X_{n}}_{X_{m}}}$$\scriptstyle{\widehat{ev}^{X_{n}}_{\mathcal{Y}}}$${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{m}}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\widehat{ev}^{X_{m}}_{\mathcal{Y}}}$${{\mathcal{P}}({\mathcal{Y}})^{X_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R^{X_{n}}_{X_{m}}}$${{\mathcal{P}}({\mathcal{Y}})^{X_{m}}.}$
Note that the commutativity of diagram (5.10) follows immediately from Equation (5.8). ∎
@@
Assume that $\mathcal{Y}$} is a Polish space and ${\mathcal{A}}_{\mathcal{Y}}$ is a countable algebra generating ${\mathcal{B}}({\mathcal{Y}})$. Then for any $\nu\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ we have
@@
there exists $\nu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ such that
@@
if and only if there exists $\mu\in{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ such that for any $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ and any $(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ we have
(5.14) $\widehat{ev}_{\mathcal{Y}}^{(m)}\big(\nu^{X_{m}}_{(A)}\big)=P_{*}\pi_{(A)}^{X_{m}}R^{\mathcal{X}}_{X_{m}}(\mu)\in{\mathcal{P}}({\Omega}_{(A)})^{X_{m}},$
equivalent, if and only if for each $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ there exists $\mu_{X_{m}}\in{\mathcal{P}}({\mathcal{Y}})^{{\mathcal{X}}_{m}}$ such that for any $(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ we have
(5.15) $\widehat{ev}_{\mathcal{Y}}^{(m)}\big(\nu^{X_{m}}_{(A)}\big)=P_{*}\pi_{(A)}^{X_{m}}(\mu_{X_{m}})\in{\mathcal{P}}({\Omega}_{(A)})^{X_{m}},$
and the system $\{R_{{\mathcal{Y}},X^{m}}^{X_{n}}:\big({\mathcal{P}}({\mathcal{Y}})^{X_{n}},\mu_{X_{n}}\big)\to\big({\mathcal{P}}({\mathcal{Y}})^{X_{m}},\mu_{X_{m}}\big),\,X_{m}\leq X_{n}\in\mathrm{P_{fin}({\mathcal{X}})}\}$ of natural projections is projective.
@@
(2) Any $\mu\in{\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$ can be written as
$\mu=\lim_{\stackrel{{\scriptstyle\leftarrow}}{{X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})}}}R^{\mathcal{X}}_{X_{m}}(\mu).$
Thus we can replace $R^{\mathcal{X}}_{X_{m}}(\mu)$ in (5.14) by $\mu_{X_{m}}$ in (5.15) in the presence of the corresponding projective system, which is the content of the last “equivalence” assertion of Theorem 5.5.
@@
(1) The equality (5.11) is a consequence of the functoriality $P_{*}:{\bf Probm}\to{\bf Meas}$ that assigns each measurable space $\mathcal{X}$} to measurable space ${\mathcal{P}}({\mathcal{X}})$ and each probabilistic morphism $T\in{\bf Probm}({\mathcal{X}},{\mathcal{Y}})$ to a measurable mapping $P_{*}T\in{\bf Meas}({\mathcal{P}}({\mathcal{X}}),{\mathcal{P}}({\mathcal{Y}}))$ [Giry82, Theorem 1], taking into account the Kolomogorov’s extension theorem.
@@
Now we assume the “if” condition. For each $X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})$, we consider the following commutative diagram:
[…]
The above half of the diagram (5.16) follows from the diagram (5.9). The lower half of the diagram (5.16) is verified straightforward.
For each $X_{n}$, we consider the projective system of probability spaces
@@
Taking into account Orbanz’s result [Orbanz2011, Theorem 1.1], see Remark 5.6(1), we conclude that there exists $\tilde{\nu}_{X_{n}}\in{\mathcal{P}}^{2}({\mathcal{Y}}^{X_{n}})$ such that

### v2

@@ 5. Probability measures on ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}$
In this Section, we assume that $\mathcal{Y}$} is a Polish space unless otherwise stated. Then ${\mathcal{P}}({\mathcal{Y}})^{k}$ and ${\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{k})$ are Polish spaces for any $k\in{\mathbb{N}}^{+}$. Using Orbanz’s description of the space ${\mathcal{P}}^{2}({\mathcal{Y}})$ for a Polish space $\mathcal{Y}$} [Orbanz2011, Theorem 1.1], we shall describe the space ${\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ using a projective system (Theorem 5.5).
@@
Recall that $R^{\mathcal{X}}_{X_{m}}$ denotes the restriction map ${\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\to{\mathcal{P}}({\mathcal{Y}})^{{\mathcal{X}}_{m}}$.
@@
Conversely, for any projective system of probability spaces $\{({\mathcal{P}}({\mathcal{Y}})^{X_{m}},\mu_{X_{m}}),R_{X_{m}}^{X_{n}}:X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})\}$ there exists a unique probability measure $\mu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ such that for all $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ we have
@@
The first assertion is straightforward. The second assertion is obtained by applying the Kolmogorov extension theorem. ∎
@@
Let ${\mathcal{A}}_{\mathcal{Y}}$ be the countable algebra generated by open balls with rational radius centered at a countable dense set in $\mathcal{Y}$}. Then ${\mathcal{A}}_{\mathcal{Y}}$ generates the Borel $\sigma$-algebra of $\mathcal{Y}$}. Set
@@
$\widehat{e}v_{\mathcal{Y}}^{(k)}:{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{k}\big)\to{\mathcal{P}}({\mathcal{Y}}^{k})$
@@
(5.2) $\widehat{ev}_{\mathcal{Y}}^{(k)}(\nu)(B_{1}\times\ldots\times B_{k}):=\int_{{\mathcal{P}}({\mathcal{Y}})^{k}}\mathfrak{m}^{k}\mu(B_{1}\times\ldots\times B_{k})\,d\nu(\mu)$
@@
$\widehat{e}v_{\mathcal{Y}}^{X_{m}}:{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{m}}\big)\to{\mathcal{P}}({\mathcal{Y}}^{X_{m}})$
@@
(5.3) $\widehat{ev}_{\mathcal{Y}}^{X_{m}}(\nu)(B_{1}\times\ldots\times B_{m}):=\int_{{\mathcal{P}}({\mathcal{Y}})^{X_{m}}}\mathfrak{m}^{m}\mu(B_{1}\times\ldots\times B_{m})\,d\nu(\mu)$
@@
(5.4) $\widehat{ev}^{(k)}_{\mathcal{Y}}=P_{*}ev_{{\mathcal{Y}}^{k}}.$
@@
(5.5) $\widehat{ev}_{\mathcal{Y}}^{\mathcal{X}}(\nu):=\lim_{\stackrel{{\scriptstyle\leftarrow}}{{X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})}}}\widehat{ev}_{\mathcal{Y}}^{X_{m}}\big((R^{\mathcal{X}}_{X_{m}})_{*}\nu\big)\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}}),$
@@
(1) Equation (5.4) follows directly from (5.2) and Equation (2.2).
(2) The second assertion follows from the first one, taking into account the measurability of $P_{*}ev_{{\mathcal{Y}}^{k}}$ (and its $(\tau_{w},\tau_{w})$-continuity if $\mathcal{Y}$} is Polish) by [Giry82, Theorem1].
3) For $X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})$ let $r^{\mathcal{X}}_{X_{m}}:{\mathcal{Y}}^{\mathcal{X}}\to{\mathcal{Y}}^{X_{m}}$ and $r^{X_{n}}_{X_{m}}:{\mathcal{Y}}^{X_{n}}\to{\mathcal{Y}}^{X_{m}}$ denote the restriction maps.
By definition of $ev_{{\mathcal{Y}}^{X_{n}}}$, we verify immediately that the following diagram
(5.6) ${{\mathcal{P}}({\mathcal{Y}})^{X_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R^{X_{n}}_{X_{m}}}$$\scriptstyle{{ev}_{{\mathcal{Y}}^{X_{n}}}}$${{\mathcal{P}}({\mathcal{Y}})^{X_{m}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{ev}_{{\mathcal{Y}}^{X_{m}}}}$${{\mathcal{Y}}^{X_{n}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{r^{X_{n}}_{X_{m}}}$${{\mathcal{Y}}^{X_{m}}}$
is commutative. Consequently, the following diagram for any $X_{m}\leq X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})$:
(5.7) ${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{n}}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}(R^{X_{n}}_{X_{m}})}$$\scriptstyle{P_{*}{ev}_{{\mathcal{Y}}^{X_{n}}}}$${{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{X_{m}}\big)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{P_{*}{ev}_{{\mathcal{Y}}^{X_{m}}}}$${{\mathcal{P}}({\mathcal{Y}}^{X_{n}})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(r^{X_{n}}_{X_{m}})_{*}}$${{\mathcal{P}}({\mathcal{Y}}^{X_{m}})}$
is commutative. Combining the commutativity of the diagram (5.7), Equation (5.4), and the second assertion of Lemma 5.4, we complete the proof of Lemma 5.4, taking into account the Kolmogorov extension theorem. ∎
@@
Assume that $\mathcal{Y}$} is a Polish space and ${\mathcal{A}}_{\mathcal{Y}}$ is the countable algebra generating ${\mathcal{B}}({\mathcal{Y}})$ defined above. Then for any $\nu\in{\mathcal{P}}({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}})$ we have
@@
there exists $\nu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ such that for any $(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ and $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ we have
@@
if and only if there exists $\mu\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$ such that for any $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ and any $(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ we have
(5.11) $\widehat{ev}_{{\Omega}_{(A)}}^{(m)}\big(\nu^{X_{m}}_{(A)}\big)=P_{*}\pi_{(A)}^{X_{m}}(r^{\mathcal{X}}_{X_{m}})_{*}\mu\in{\mathcal{P}}({\Omega}_{(A)}^{X_{m}}).$
Equivalent, there exists $\nu\in{\mathcal{P}}\big({\mathcal{P}}({\mathcal{Y}})^{\mathcal{X}}\big)$ such that (5.10) holds, if and only if for each $X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})$ there exists $\mu_{X_{m}}\in{\mathcal{P}}({\mathcal{Y}}^{{\mathcal{X}}_{m}})$ such that for any $(A)\in{\mathcal{H}}({\mathcal{A}}_{\mathcal{Y}})$ we have
(5.12) $\widehat{ev}_{{\Omega}_{(A)}}^{(m)}\big(\nu^{X_{m}}_{(A)}\big)=P_{*}\pi_{(A)}^{X_{m}}(\mu_{X_{m}})\in{\mathcal{P}}({\Omega}_{(A)}^{X_{m}}),$
where the system $\{r_{X^{m}}^{X_{n}}:({\mathcal{Y}}^{X_{n}},\mu_{X_{n}})\to({\mathcal{Y}}^{X_{m}},\mu_{X_{m}}),\,X_{m}\leq X_{n}\in\mathrm{P_{fin}({\mathcal{X}})}\}$ of probability spaces is projective.
@@
(2) Any $\mu\in{\mathcal{P}}({\mathcal{Y}}^{\mathcal{X}})$ can be written as
$\mu=\lim_{\stackrel{{\scriptstyle\leftarrow}}{{X_{m}\in\mathrm{P_{fin}}({\mathcal{X}})}}}(r^{\mathcal{X}}_{X_{m}})_{*}(\mu).$
By the Kolmogorov extension theorem, we can replace $(r^{\mathcal{X}}_{X_{m}})_{*}\mu$ in (5.11) by $\mu_{X_{m}}$ in (5.12) in the presence of the corresponding projective system, which is the content of the last “equivalence” assertion of Theorem 5.5.
@@
(1) The equality (5.8) is a consequence of the functoriality $P_{*}:{\bf Probm}\to{\bf Meas}$ that assigns each measurable space $\mathcal{X}$} to measurable space ${\mathcal{P}}({\mathcal{X}})$ and each probabilistic morphism $T\in{\bf Probm}({\mathcal{X}},{\mathcal{Y}})$ to a measurable mapping $P_{*}T\in{\bf Meas}({\mathcal{P}}({\mathcal{X}}),{\mathcal{P}}({\mathcal{Y}}))$ [Giry82, Theorem 1].
@@
Now we assume the “if” condition of Theorem 5.5. For each $X_{n}\in\mathrm{P_{fin}}({\mathcal{X}})$, we consider the projective system of probability spaces
@@
Taking into account Orbanz’s result [Orbanz2011, Theorem 1.1] for ${\mathcal{P}}^{2}({\mathcal{Y}}^{X_{n}})$, see Remark 5.6(1), we conclude that there exists $\tilde{\nu}_{X_{n}}\in{\mathcal{P}}^{2}({\mathcal{Y}}^{X_{n}})$ such that